40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1930     Entry time: Wed Aug 19 23:57:35 2009
Author: rob 
Type: Update 
Category: Locking 
Subject: report 

 

locking work proceeding apace tonight.

diagonalized DRM with setDDphases & senseDRM

initial locks are fairly quick, aqstep script succeeds reliably.

first part of cm_step (handoff CARM-> MCL) usually works.

tuning up later parts of cm_step (presumably due to optical gain changes resulting from MOPA decline). 

got to arm powers ~60.

ELOG V3.1.3-